jigglypuff314 (jigglypuff314):

Part A: Find the point on the parabola y^2 = 2x that is closest to point (1, 4) which I got to be (2, root(5) ) then I need help with PART B: Verify that the tangent to the parabola y^2 = 2x at your answer is perpendicular to the line connecting your answer to the point (1, 4) could someone just explain to me the process of Part B?

OpenStudy (anonymous):

let (x,y)be the point close to (1,4) then distance \[d=\sqrt{\left( x-1 \right)^{2}+\left( y-4 \right)^{2}}\] now (x1,y1) lies on parabola \[(y)^{2}=2x,x=\frac{ y ^{2} }{ 2 }\] \[d=\sqrt{\left( \frac{ y ^{2} }{ 2 }-1 \right)^{2}+\left( y-4 \right)^{2}}\] Let \[D=d ^{2}=\left( \frac{ y ^{2} }{ 2 }-1 \right)^{2}+\left( y-4 \right)^{2}\] \[\frac{ dD }{ dy }=2\left( \frac{ y ^{2} }{ 2 } -1\right)\frac{ 2y }{ 2 }+2(y-4)\] \[=y ^{3}-2y+2y-8\] \[=y ^{3}-8\] \[\frac{ dD }{ dy }=0 gives y ^{3}-8=0,y=2\] \[\frac{ d ^{2}D }{ dy ^{2} }=3y ^{2}\] \[at~y=2,\frac{ d ^{2}D }{ dy ^{2} }=3*2^{2}=12>0\] hence D is minimum at y=2 \[2x=y ^{2}=2^{2}=4,x=2\] Hence point nearest to (1,4) is (2,2)

OpenStudy (anonymous):

find dy/dx at (2,2) from y^2=2x this is the slope of tangent also find the slope of line through (1,4) and (2,2) multiply them . if product is-1 then they are perpendicular.

OpenStudy (anonymous):

correction above write only (x,y) in place of (x1,y1)